# Small Drafts, Big Verdict: Information-Intensive Visual Reasoning via Speculation

### v1

@@ 4.4 Ablation Study
Impact of Verdict Stage. The verdict stage yields higher performance than majority voting across information-intensive benchmarks (Figure 8). Notably, majority voting with all five draft experts performs comparably as majority voting with three draft experts, consistent with our finding that consensus selection can match the performance of all drafts at a lower cost (Figure 6). SV further surpasses both by leveraging the verdict’s error correction ability, successfully capturing minority-correct cases that majority voting discards (Figure 4 and Figure 3).
@@
Choice of Verdict Scale. Using a large verdict model yields stronger gains than a small verdict model. For ablations, we select GLM-4.1V-9B-Thinking as the small verdict because it is the strongest reasoning model among the baselines. However, results in Table 3 show that it brings only modest improvements, while GPT-4o delivers additional gains of 3.4% on InfographicVQA and 1.3% on ChartMuseum compared to this small verdict. These results indicate that even reasoning-strong small verdicts offer limited benefit in synthesizing correct answers, validating SV’s design principle of invoking a strong verdict only once to achieve robust and efficient error correction.
@@
Figure 8: Performance comparison on SV and majority voting with different model sets.
@@
Reasoning paths 88.4 64.0
Table 3: Ablations on verdict scale. A subset of 1000 samples is tested on InfographicVQA.
@@
Reasoning baseline 84.5 48.0 57.3
GLM-4.1V-9B-Thinking Verdict 86.0 48.0 59.4
GPT-4o Verdict 89.4 49.3 64.0
@@ 6 Ethics Statement
This work does not involve human subjects, sensitive personal data, biometrics, or medical information. All datasets used are publicly available under permissible licenses and are not privacy-sensitive. We recognize that any automated reasoning system may produce incorrect or misleading outputs. To ensure responsible use, we emphasize that our method is intended for research and analysis rather than deployment in high-stakes settings. Users are encouraged to verify model outputs and apply human oversight when necessary. We take full responsibility for all reported results, analyses, and claims, and we welcome community scrutiny and feedback.
@@ 7 Reproducibility Statement
To support reproducibility, we provide comprehensive implementation details throughout our paper. Key experimental configurations, such as draft expert selection, consensus scoring computation, and verdict model specifications, are documented in Section 3.4 and Section 4.1. Detailed prompt templates are presented in Appendix H. The code is released to further clarify the implementation steps and enable faithful reproduction of our results.
@@ Appendix A Dataset statistics
Table 4 reports the statistics of the four evaluation benchmarks. All benchmarks are based on real-world images rather than synthetic renderings, ensuring the authenticity and diversity of the evaluation setting. In particular, InfographicVQA, ChartMuseum, and ChartQAPro are information-intensive benchmarks: they contain thousands of images and questions with dense textual and numerical content, collected from diverse sources spanning 2594, 157, and 184 distinct web domains respectively (Mathew et al., 2021; Tang et al., 2025; Masry et al., 2025). This diversity reduces source bias and reflects practical challenges in multimodal reasoning.
HR-Bench 4K is used primarily to evaluate the generalization of our method, serving as a high-resolution benchmark with average sizes exceeding 4000×3500 pixels (Wang et al., 2025b). At the same time, one of our main benchmarks, InfographicVQA, also exhibits high-resolution characteristics. In particular, it frequently contains long-format images where diagrams span large vertical layouts (see the case in Figure 3), which further compounds the difficulty of grounding and multi-hop reasoning across dispersed regions.
@@
Appendix B Costs
@@
Table 5: Average inference cost of GPT-4o as verdict per sample across benchmarks. Costs are computed using GPT-4o (gpt-4o-2024-08-06) pricing by September 2025.
@@
Appendix C Supplementary recovery analysis on information-intensive benchmarks
Table 6 and Figure 9 show the detailed recovery statistics across information-intensive benchmarks with GPT-4o as verdict. We break down SV’s performance by expert correctness: (i) cases where the majority of draft experts are correct (majority-correct), (ii) cases where only a minority are correct (minority-correct), (iii) cases where none are correct (zero-correct). While the main paper focuses on the GPT-4o’s error cases to isolate SV’s effectiveness, we provide the full results here for completeness.
Notably, in the zero-correct setting, recovery occurs rarely (2.6-24%), but it demonstrates verdict’s surprising ability to infer the correct answer by synthesizing signal from entirely noisy reasoning.
@@ Appendix C Supplementary recovery analysis on information-intensive benchmarks
Figure 9: SV’s correction ability on verdict’s correct cases (GPT-4o as verdict), complementary to its error cases in the Figure 4.
Appendix D Ablation study on model pool compositions
Evaluation with 7-9B Model Pool (Non-Thinking). SV maintains its effectiveness when replacing thinking models with faster non-thinking alternatives. Specifically, we replace the two thinking models in our original pool (i.e., GLM-4.1V-9B-Thinking (Team et al., 2025b) and MiMO-VL-7B-RL (Xiaomi, 2025)) with non-thinking models (i.e., LLaVA-OneVision-1.5-8B (An et al., 2025), and Eagle 2.5-8B (Chen et al., 2025)), while keeping the remaining three models unchanged. While these substitutes sacrifice some reasoning capability, they enable faster inference. As shown in Table 7, with GPT-4o as verdict, SV achieves 86.3% on InfographicVQA under this configuration, surpassing all baselines. Notably, SV outperforms the best draft expert by 4.6% and exceeds the large 72B model by 1.9%. These results demonstrate that SV achieves strong performance by integrating reasoning paths from individually weaker but faster models.
Evaluation with 2-4B Model Pool. We also evaluate SV on an even smaller model pool consisting of 2-4B models: Qwen2.5-VL-Instruct-3B (Bai et al., 2025), LLaVA-OneVision-1.5-4B (An et al., 2025), InternVL3.5-4B (Wang et al., 2025a), Gemma 3-4B (Team et al., 2025a), and Ovis2.5-2B (Lu et al., 2025). As shown in the Table 8, with GPT-4o as verdict, SV achieves 84.5% on InfographicVQA, surpassing the best draft expert by 9.5% and the 72B baseline by 0.3%. This demonstrates SV’s ability to extract effective collective reasoning even from significantly weaker individual models, confirming the robustness of our paradigm across varying model scales.
Table 7: Results on test sets of three information-intensive benchmarks with 7-9B draft experts. We compare SV against closed-source VLMs, open-source VLMs, and the tool-driven method, with all results reproduced by ourselves. The best results for each benchmark are highlighted in bold and the second-best results are underlined.
@@ Appendix D Ablation study on model pool compositions
SV w/ Qwen2.5-VL-72B-Instruct Verdict 84.2
Table 8: Results on test sets of three information-intensive benchmarks with 2-4B draft experts. Same evaluation and marking conventions as Table 7.
@@
SV w/ Qwen2.5-VL-72B-Instruct Verdict 80.4
Appendix E Ablation studies on verdict input configuration
E.1 Impact of visual input to verdict

### v2

@@ 4.3 Results on High-Resolution Benchmark
4.4 Results on Multimodal Reasoning Benchmarks
To examine SV’s broader generalization, we further evaluate it on two additional multimodal reasoning benchmarks: TallyQA (counting) and MathVista (mathematical reasoning). Table 2 shows that SV provides consistent gains, outperforming the strongest draft expert by 3.0% on TallyQA-Complex and 2.6% on MathVista, and improving over GPT-4o by 1.5% and 17.8%, respectively. These results demonstrate that SV generalizes to diverse visual reasoning tasks.
@@ 4.5 Ablation Study
Impact of Verdict Stage. The verdict stage yields higher performance than majority voting across information-intensive benchmarks (Figure 8). Notably, majority voting with all five candidate models performs comparably as majority voting with three draft experts, consistent with our finding that consensus selection can match the performance of all drafts at a lower cost (Table 6). SV further surpasses both by leveraging the verdict’s error correction ability, successfully capturing minority-correct cases that majority voting discards (Figure 4 and Figure 3).
Beyond majority voting, SV consistently outperforms a strong LMM-as-a-Judge baseline (LLaVA-Critic-72B (llavacritic)) by 4.9–11.9% (Figure 8). In this setting, LLaVA-Critic scores or ranks the same draft experts selected by SV and outputs the single best candidate. This advantage is attributed to SV’s synthesis-based verdict, which cross-checks and integrates complementary factual cues across multiple trajectories, rather than relying on selecting one trajectory that may be favored by reasoning style. Detailed ablations with LLaVA-Critic are provided in Appendix G.
@@
Choice of Verdict Scale. Using a large verdict model yields stronger gains than a small verdict model. We evaluate three strong small verdicts (i.e., GLM-4.1V-9B-Thinking, MiMO-VL-7B-RL, Qwen2.5-VL-7B-Instruct), and all of them underperform SV across benchmarks (Table 4). Notably, the small reasoning verdicts generate 60-200x more output tokens yet still yield weaker performance, indicating worse cost-efficiency trade-offs (details in Appendix C). The results validate SV’s design principle of invoking a strong verdict only once to achieve robust and efficient error correction.
@@
Figure 8: Performance comparison on SV, majority voting and LLaVA-Critic with different model sets.
@@
Reasoning paths (SV) 88.4 64.0
Table 4: Ablations on verdict scale.
@@
Reasoning baseline 84.8 48.0 57.3
GLM-4.1V-9B-Thinking Verdict 84.7 48.0 59.4
MiMO-VL-RL-7B Verdict 85.4 46.9 60.3
Qwen2.5-VL-Instruct 7B Verdict 84.1 47.1 57.2
GPT-4o Verdict 88.4 49.3 64.0
4.6 Cost-Efficiency Analysis
We quantify SV’s cost-efficiency by comparing it against large reasoning model baselines, including GPT-4o and o1. As described in Section 3.3, SV achieves cost-efficiency through its verdict stage by concentrating computation in prefill and substantially reducing expensive autoregressive decoding.
As shown in Table 5 and 6, SV consistently improves over GPT-4o by 6.6–12.0% across datasets while maintaining comparable cost. Compared with o1, SV attains substantially better performance on InfographicVQA and ChartQAPro and comparable performance on ChartMuseum while requiring only 15–26% of o1’s cost. These results confirm that SV delivers markedly superior cost-efficiency, achieving and even outperforming o1-level performance with lower computational cost.
Table 5: Performance on 1000 randomly sampled test instances per benchmark.
Method InfographicVQA ChartMuseum ChartQAPro
GPT-4o 76.3 42.7 51.7
GPT-o1 77.8 50.6 58.8
SV w/ GPT-4o Verdict 88.3 49.3 63.4
Table 6: Average inference API cost per sample across benchmarks.
Method InfographicVQA ChartMuseum ChartQAPro
GPT-4o \$0.0038 \$0.0213 \$0.0210
GPT-o1 \$0.0263 \$0.0663 \$0.0478
SV w/ GPT-4o Verdict \$0.0068 \$0.0109 \$0.0071
@@ 6 Ethics Statement
All authors have read and commit to adhering to the ICLR Code of Ethics. This work does not involve human subjects, sensitive personal data, biometrics, or medical information. All datasets used are publicly available under permissible licenses and are not privacy-sensitive. We recognize that any automated reasoning system may produce incorrect or misleading outputs. To ensure responsible use, we emphasize that our method is intended for research and analysis rather than deployment in high-stakes settings. Users are encouraged to verify model outputs and apply human oversight when necessary. We take full responsibility for all reported results, analyses, and claims, and we welcome community scrutiny and feedback.
@@ 7 Reproducibility Statement
To support reproducibility, we provide comprehensive implementation details throughout our paper. Key experimental configurations, such as draft expert selection, consensus scoring computation, and verdict model specifications, are documented in Section 3.4 and Section 4.1. Detailed prompt templates are presented in Appendix M. The supplementary material includes anonymized source code to further clarify the implementation steps and enable faithful reproduction of our results.
@@ Appendix A Dataset Statistics
Table 7 reports the statistics of the four major evaluation benchmarks. All benchmarks are based on real-world images rather than synthetic renderings, ensuring the authenticity and diversity of the evaluation setting. In particular, InfographicVQA, ChartMuseum, and ChartQAPro are information-intensive benchmarks: they contain thousands of images and questions with dense textual and numerical content, collected from diverse sources spanning 2594, 157, and 184 distinct web domains respectively (mathew2021infographicvqa; tang2025chartmuseum; masry2025chartqapro). This diversity reduces source bias and reflects practical challenges in multimodal reasoning.
HR-Bench 4K is used primarily to evaluate the generalization of SV, serving as a high-resolution benchmark with average sizes exceeding 4000×3500 (wang2025divide). Meanwhile, one of our main benchmarks, InfographicVQA, also exhibits high-resolution characteristics. In particular, it contains many images where diagrams span large vertical layouts (see the case in Figure 3), which further compounds the difficulty of grounding and multi-hop reasoning across dispersed regions.
@@
Appendix B Additional Related Work
Large Language Model Ensemble. Majority voting aggregates answers by frequency, but fails when the correct solution is produced by a minority. Universal Self-Consistency (chen2023universal) mitigates this failure mode by prompting the LLM to select the most consistent candidate across samples. Further, learned aggregators read multiple rationales and synthesize them to recover minority-correct information (qi2025learning; zhao2025majority). However, these approaches focus on text-only ensembling. In vision-language reasoning, supervision of ensembling is not cost-effective since multimodal complexity requires costly, fine-grained annotations.
Appendix C Additional Cost–Efficiency Analysis
C.1 Average API Cost of SV with GPT-4o Verdict
@@
Table 8: Average inference cost of GPT-4o as verdict per sample across benchmarks. Costs are computed using GPT-4o (gpt-4o-2024-08-06) pricing by November 2025.
@@
C.2 Average Output Tokens across Verdicts
To compare the token efficiency of replacing the large verdict with smaller models, we keep the draft stage identical and substitute the verdict with different draft models. All verdict models are prompted to output only the final answer, and their performances are reported in Table 4 in the main paper.
As shown in Table 9, the small reasoning models, GLM-4.1V-9B-Thinking and MiMO-VL-7B-RL, generate 60–200× more output tokens than GPT-4o, yet still underperform SV. This highlights SV’s advantage: using a strong verdict once for compact verification and synthesis is substantially more token-efficient than relying on smaller verdicts that require long autoregressive reasoning to reach a decision.
Table 9: Average verdict output tokens per sample across benchmarks under different verdicts.
Verdict model InfographicVQA ChartMuseum ChartQAPro
GLM-4.1V-9B-Thinking 272.1 604.5 440.5
MiMO-VL-7B-RL 183.2 368.2 378.1
Qwen2.5-VL-7B-Instruct 3.3 3.4 2.8
GPT-4o (SV) 2.7 3.0 2.2
Appendix D Supplementary Recovery Analysis on Information-Intensive Benchmarks
Table 10 and Figure 9 show the detailed recovery statistics across benchmarks with GPT-4o as verdict. We break down SV’s performance by expert correctness: (i) cases where the majority of draft experts are correct (majority-correct), (ii) cases where only a minority are correct (minority-correct), (iii) cases where none are correct (zero-correct). While the main paper focuses on the GPT-4o’s error cases to isolate SV’s effectiveness, we provide the full results here for completeness. Notably, in the zero-correct setting, recovery occurs rarely (2.6-24%), but it demonstrates verdict’s surprising ability to infer the correct answer by synthesizing signal from entirely noisy reasoning.
@@ Appendix D Supplementary Recovery Analysis on Information-Intensive Benchmarks
Figure 9: SV’s correction ability on verdict’s correct cases (GPT-4o as verdict), complementary to its error cases in Figure 4.
Table 11: Results on test sets of InfographicVQA with 7–9B draft experts. We compare SV against closed-source VLMs, open-source VLMs, and tool-driven methods. $\dagger$ denotes results reported in the original papers; all other results are reproduced by us. The best results for each benchmark are highlighted in bold and the second-best results are underlined.
@@
  $\Delta$ (vs. GPT-4o) +9.8
Table 12: Results on test sets on InfographicVQA with 2–4B draft experts. Same evaluation and marking conventions as Table 11.
@@
  $\Delta$ (vs. GPT-4o) +8.0
Appendix E Ablation Study on Model Pool Compositions
Evaluation with 7-9B Model Pool (Non-reasoning). SV maintains its effectiveness when replacing reasoning models with faster non-reasoning alternatives. Specifically, we replace the two reasoning models in our original pool (i.e., GLM-4.1V-9B-Thinking (glm) and MiMO-VL-7B-RL (Yue2025MiMo)) with non-reasoning models (i.e., LLaVA-OneVision-1.5-8B (2025llava), and Eagle 2.5-8B (2025eagle)), while keeping the remaining three models unchanged. While these substitutes sacrifice some reasoning capability, they enable faster inference. As shown in Table 11, with GPT-4o as verdict, SV achieves 86.3% on InfographicVQA under this configuration, surpassing all baselines. Notably, SV surpasses the best draft expert by 4.6% and improves over GPT-4o by 9.8%. These results demonstrate that SV achieves strong performance by integrating reasoning paths from individually weaker but faster models.
Evaluation with 2-4B Model Pool. We also evaluate SV on an even smaller model pool consisting of 2-4B models: Qwen2.5-VL-3B-Instruct (bai2025qwen2), LLaVA-OneVision-1.5-4B (2025llava), InternVL3.5-4B (2025internvl3), Gemma 3-4B (2025gemma), and Ovis2.5-2B (lu2025ovis2). As shown in the Table 12, with GPT-4o as verdict, SV achieves 84.5% on InfographicVQA, surpassing the best draft expert by 9.5% and yielding an 8% gain over GPT-4o. This demonstrates SV’s ability to extract effective collective reasoning even from significantly weaker individual models, confirming the robustness of our paradigm across varying model scales.
Appendix F Analysis of Consensus Selection Mechanism
This section provides a detailed analysis of our consensus expert selection mechanism, addressing three aspects: (1) selection frequency distribution, (2) impact of normalization on consensus scores, and (3) impact of NLL estimation strategy.
F.1 Selection Frequency Distribution
To understand how the consensus mechanism selects draft experts, we analyze the selection frequency of each model across benchmarks. Table 13 reports the proportion of instances in which each model is selected. All models participate with non-trivial frequencies (38.6%–84.7%), with no model dominating or being marginalized. This indicates that different drafts specialize in different subsets of questions, and the consensus mechanism leverages complementary agreement and disagreement across experts rather than collapsing to a single always-selected model.
Table 13: Selection frequency of each draft model across benchmarks.
Model InfographicVQA ChartQAPro
Qwen2.5-VL-7B-Instruct 84.7% 77.7%
GLM-4.1V-9B-Thinking 38.6% 69.3%
MiMO-VL-7B-RL 54.8% 56.8%
InternVL3-8B 73.4% 53.1%
Ovis2.5-9B 48.5% 43.1%
F.2 Impact of Normalization on Consensus Scores
We further ablate the effect of normalization on consensus scores, which is motivated by the calibration gap across draft experts mentioned in Section 3.4. Different VLMs produce perplexity scores on systematically different numerical scales due to training and tokenization differences. In our pool, for example, Qwen2.5-VL and GLM-4.1V-Thinking tend to output larger-magnitude NLLs. Without normalization, such scale mismatch causes high-magnitude models to dominate the consensus even when they do not genuinely agree more often, thereby reducing selection diversity.
We evaluate normalization on two draft pools on InfographicVQA: (i) the main pool used in our primary experiments, mixing reasoning and non-reasoning drafts, and (ii) an additional pool consisting only of non-reasoning drafts, introduced in Appendix E. Tables 14(a) and 14(c) show that, on the main pool, removing normalization collapses selection to a near-fixed subset: Qwen and GLM are selected in almost all instances. This collapse hurts performance on ChartQAPro, demonstrating that calibration artifacts can actively degrade consensus quality. Although InfographicVQA remains similar without normalization, this results from an accidental strong pairing being repeatedly selected rather than a principled aggregation.
To verify that this trend is not pool-specific, we repeat the ablation on the additional non-reasoning pool. Tables 14(b) and 14(d) show the same selection-collapse pattern without normalization, together with a clear performance drop. Across both pools, normalization mitigates calibration bias, prevents dominance by a few drafts, and yields more stable performance.
Table 14: Normalization ablations on the main and additional draft pools.
(a) Ablations on normalization on the main draft pool.
Variant InfographicVQA ChartQAPro
SV w/o norm. 88.9 59.4
SV 88.4 64.0
(b) Ablations on normalization on the additional non-reasoning pool.
Variant InfographicVQA
SV w/o norm. 84.6
SV 86.3
(c) Selection frequency on the main pool without normalization.
Model InfographicVQA ChartQAPro
Qwen2.5-VL-7B-Instruct 99.9% 100.0%
GLM-4.1V-9B-Thinking 77.1% 99.9%
MiMO-VL-7B-RL 57.0% 71.5%
InternVL3-8B 47.0% 21.4%
Ovis2.5-9B 19.1% 2.2%
(d) Selection frequency on the additional non-reasoning pool.
Model w/ norm. w/o norm.
Qwen2.5-VL-7B-Instruct 87.3% 99.9%
Eagle2.5-8B 65.4% 98.0%
LLaVA-OneVision-1.5-8B 28.6% 79.3%
InternVL3-8B 74.5% 16.7%
Ovis2.5-9B 44.2% 6.11%
F.3 Impact of NLL Estimation Strategy
We examine whether computing NLL only on final answers introduces off-policy bias, since draft experts’ answers are generated together with reasoning trajectories. Table 15 compares expert selection using answer-only NLL (SV) versus full-trajectory NLL. The two variants achieve very similar performance, suggesting that any off-policy bias from answer-only scoring is negligible for this task. Moreover, answer-only scoring is more computationally efficient, as it avoids computing NLL over long reasoning traces with many extra tokens; answer-only NLL also provides a cleaner signal by avoiding noise from diverse reasoning styles across models. We therefore use answer-only NLL in the main experiments.
Table 15: Ablation on NLL estimation strategy.
Scoring variant InfographicVQA ChartQAPro
Answer+reasoning NLL 87.9 64.3
Answer-only NLL (SV) 88.4 64.0
Appendix G Comparison to LLaVA-Critic (LMM-as-a-Judge Baseline)
We compare SV to a learned LMM-as-a-Judge baseline, LLaVA-Critic (llavacritic). LLaVA-Critic is trained as a generalist multimodal evaluator that jointly leverages visual evidence and textual reasoning to score or rank multiple candidates.
Experimental setup. We follow the official LLaVA-Critic evaluation template and consider two judging modes: (i) pointwise scoring, where the judge assigns a scalar score to each candidate and selects the highest-scored one; (ii) pairwise ranking, where candidates are compared (i.e., first two candidates are compared, and the winner is then compared against the third in our scenario). To ensure a comprehensive comparison, we evaluate LLaVA-Critic under two settings: (1) judging the same three consensus experts selected by SV (directly replacing the verdict); (2) judging five candidate models. In all cases, the judge outputs a single best candidate.
For SV, we conduct experiments with two verdict models: GPT-4o and LLaVA-OneVision-7B (li2024llava). GPT-4o serves as a strong proprietary verdict model, matching our main experiments, while LLaVA-OneVision-7B shares the same backbone as LLaVA-Critic, enabling a fairer comparison in the open-source regime.
Results. Table 16 shows that when judging the same three draft experts, SV consistently outperforms LLaVA-Critic-7B across all three benchmarks under both verdict choices. With GPT-4o as the verdict, SV improves over LLaVA-Critic-7B by 4.9–11.9%. With LLaVA-OneVision-7B as the verdict, SV still yields gains of 0.5–6.6%, which demonstrates that the advantage stems from SV’s design rather than from verdict model superiority alone.
This gap is attributed to the different aggregation principles: LLaVA-Critic performs selection-based judging and can only pick one trajectory, while SV performs synthesis-based verification, cross-checking and integrating complementary factual cues across trajectories. In practice, the critic is also more sensitive to surface form (e.g., verbosity, repetitiveness, or stylistic fluency), whereas SV is guided by consensus evidence and is therefore more robust to such artifacts.
Table 17 further reports the total judge/verdict tokens. Under comparable prefill-based budgets, SV achieves higher performance with similar or lower token usage on InfographicVQA and ChartMuseum, and still attains a much better cost–performance trade-off on ChartQAPro.
When increasing the candidate pool to five models (Table 18), LLaVA-Critic-7B remains clearly below SV, indicating that simply adding more candidates to a selection-based judge does not close the gap. Finally, even a much larger judge (LLaVA-Critic-72B, Table 19) is still notably weaker than SV, even including the variant with the smaller LLaVA-OneVision-7B verdict. This confirms that SV is more effective than selection-based judging for information-intensive visual reasoning.
Table 16: Performance of SV and LLaVA-Critic-7B.
Method InfographicVQA ChartMuseum ChartQAPro
LLaVA-Critic-7B (pointwise) 83.5 40.1 52.4
LLaVA-Critic-7B (pairwise) 81.4 38.9 52.1
SV w/ GPT-4o verdict 88.4 49.3 64.0
SV w/ LLaVA-OneVision-7B verdict 84.0 44.1 59.0
Table 17: Average judge/verdict tokens per sample.
Method InfographicVQA ChartMuseum ChartQAPro
LLaVA-Critic-7B (pointwise) 1053.3 3723.7 1290.7
LLaVA-Critic-7B (pairwise) 1342.9 4689.4 1759.1
SV w/ GPT-4o verdict 701.7 3122.3 1441.7
SV w/ LLaVA-OneVision-7B verdict 702.4 3123.0 1442.4
Table 18: Performance of SV and LLaVA-Critic-7B judging five candidate models (1k samples).
Method InfographicVQA ChartMuseum ChartQAPro
LLaVA-Critic-7B (pointwise) 82.3 34.6 56.5
SV w/ GPT-4o verdict 88.3 49.3 63.4
SV w/ LLaVA-OneVision-7B verdict 83.6 44.1 57.8
Table 19: Performance of SV and LLaVA-Critic-72B judging three consensus experts (1k samples).
Method InfographicVQA ChartMuseum ChartQAPro
LLaVA-Critic-72B (pointwise) 82.6 40.4 53.5
LLaVA-Critic-72B (pairwise) 83.1 43.0 55.8
SV w/ GPT-4o verdict 88.3 49.3 63.4
SV w/ LLaVA-OneVision-7B verdict 83.6 44.1 57.8
Appendix H Ablation Studies on Verdict Input Configuration
H.1 Impact of Visual Input to Verdict